Consider a project to supply 102 million postage stamps per year to the U.S. Postal Service for the next five years. You have an idle parcel of land available that cost $1,920,000 five years ago; if the land were sold today, it would net you $2,120,000 aftertax. The land can be sold for $2,320,000 after taxes in five years You will need to install $5.42 million in new manufacturing plant and equipment to actually produce the stamps; this plant and equipment will be depreciated straight-line to zero over the projects five-year life The equipment can be sold for $520,000 at the end of the project. You will also need $620,000 in initial net working capital for the project, and an additional investment of $52,000 in every year thereafter. Your production costs are 0.52 cents per stamp, and you have fixed costs of $1,070,000 per year. If your tax rate is 30 percent and your required return on this project is 10 percent, what bid price should you submit on the contract? (Do not round intermediate calculations and round your final answer to 5 decimal places. (e.g., 32.16161)) Bid price
Answer #1

Bid price per stamp : $ 0.54864

Workings:

Initial investment: $ 8,160,000

Installed cost of equipment $ 5,420,000
Initial working capital 620,000
Current cost of land ( opportunity cost, relevant) 2,120,000
Purchase cost of land ( sunk cost, hence not relevant) 0
Total Initial Investment $ 8,160,000

Terminal value : $ 3,512,000

After-tax sale proceeds of land 2,320,000
After-tax sale proceeds of equipment 364,000
Working capital recapture 828,000
Total terminal cash flows 3,512,000

Operating Cash flows after taxes = EBITDA x ( 1 - t ) + Depreciation x t = [ 102,000,000 P - 53,040,000 - 1,070,000 ] x ( 0.70) + 1,084,000 x 0.30 = 71,400,000 P - 37,877,000 + 325,200 = 71,400,000 P - 37,551,800

Net cash flows ( Year 1 - Year 4) = 71,400,000 P - 37,551,800 - 52,000 = 71,400,000 P - 37,603,800

Total cash flow ( Year 5 ) = 71,400,000 P - 37,551,800 + 3,512,000 = 71,400,000 P - 34,039,800

Present value of cash inflows ( Years 1 - Year 4 ) = ( 71,400,000 P - 37,603,800 ) x 3.1699 = 226,330,860 P - 119,200,286

Present value of cash inflows ( Year 1 ) = ( 71,400,000 P - 34,039,800 ) x 0.6209 = 44,332,260 P - 21,135,312

At financial break-even, NPV = 0

270,663,120 P - 140,335,598 = 8,160,000

or P = $ 0.54864